# Robust Feedback Optimization with Model Uncertainty: A Regularization Approach

### v1

@@ V Numerical Experiments
where $R\in\mathbb{R}^{2n\times 2n}$ and $Q\in\mathbb{R}^{n\times n}$ are positive definite cost matrices, $\mathcal{U}_{i,k}\triangleq\{[p_{i},\,q_{i}]:0\leq p_{i}\leq p^{\text{MPP}}_{i,k},q_{\min}\leq q_{i}\leq q_{\max}\}$ is the constraint set, and $q_{\min}$ and $q_{\max}$ are lower and upper bounds on reactive actuation, respectively.
We adopt the UNICORN 56-bus test case [35] with 5 photovoltaic inverters. Although the input-output sensitivity is a nonlinear function of $u_{k}$, we learn a constant approximation $\hat{H}$ based on power flow linearization and historical data of the injected powers and voltages. This sensitivity becomes even more inexact when the distribution grid experiences a topology change, specifically when the point of common coupling is switched from bus 1 to bus 26. While the uncertainty set for $\hat{H}$ is hard to characterize correctly, we tune the regularizers of (13) and (14) by gradually decreasing their values from conservative upper bounds. We augment the standard feedback optimization controller (4) and the proposed controllers (13) and (14) with projection to $\mathcal{U}_{i,k}$, use the same step size for these controller, and apply them to the changed grid.
As shown in the first sub-figure of Fig. 2(a), when implemented in a new environment with sensitivity uncertainty, standard feedback optimization causes oscillations and voltage violation. Note that the dashed lines in the sub-figures on the first row denote the maximum and minimum voltage limits, which equal $1.1$ p.u. and $0.9$ p.u., respectively. This standard controller also requires large reactive power actuation. In contrast, robust feedback optimization controllers maintain voltage stability after the point of common coupling changes. This is achieved by conservatively regulating control inputs, a consequence of regularization in the face of uncertainty. Furthermore, as shown in the sub-figures on the last row, the proposed controllers lead to less active power curtailments compared to the standard approach. Overall, robust feedback optimization effectively handles model uncertainty in this example of real-time voltage regulation.
@@ VI Conclusion
We addressed steady-state optimization of a dynamical system subject to model uncertainty by presenting robust feedback optimization, which seeks optimal closed-loop performance given all possible sensitivities falling within bounded uncertainty sets. Tractable reformulations for this min-max problem via regularized steady-state specifications were then provided. We also showcased the adaptation and robustness of our controllers through theoretical tracking guarantees and numerical experiments on a distribution grid. Future avenues include tuning regularizers via differentiable programming, incorporating regularization into online sensitivity learning, and pursuing robustness against model uncertainty for nonlinear stable systems.
@@
We provide two lemmas that quantify the state dynamics (1) and the controller iterations (13) and (14).
@@
Let the conditions of Theorem 2 hold. The state dynamics (1) ensure
@@
We start by analyzing the dynamics (1). Recall that $x_{\textup{ss},k}=(I-A)^{-1}(Bu_{k}+d_{x,k})$ is the steady state of (1) induced by $u_{k}$ and $d_{x,k}$. Let $L_{x}^{u}\triangleq\|(I-A)^{-1}B\|$ and $L_{x}^{d}\triangleq\|(I-A)^{-1}\|$ be the Lipschitz constants of $x_{\textup{ss},k}$ with respect to $u$ and $d$, respectively. Since $\rho(A)<1$, for any given positive definite $\bar{Q}\in\mathbb{R}^{n\times n}$, there exists a unique positive definite $P\in\mathbb{R}^{n\times n}$ satisfying the Lyapunov equation $A^{\top}PA-P+\bar{Q}=0$. Therefore,
@@
^$\hat{\nabla}\Phi_{\ell_{2},k}(u_{k},y_{k})$ $=2Ru_{k}\!+\!2\lambda\hat{H}^{\top}Q(y_{k}\!-\!r_{k})\!+\!2\rho_{\text{gen}}u_{k},$
@@
$\lVert u_{k+1}-u_{k}\rVert=\|u_{k+1}\!-\!\eta\hat{\nabla}\Phi_{\ell_{2},k}(u_{k},y_{k})\!-\!u_{k}\|$
@@
where (a) uses the triangle inequality; (b) is because of the triangle inequality and $u_{k}^{*}$ being the optimal point, i.e., $\nabla\Phi_{\ell_{2},k}(u_{k}^{*})=0$; (c) follows from the Lipschitz continuity of $\nabla\Phi_{\ell_{2},k}$, the addition and subtraction of $Hu_{k}+d_{k}$ inside $\|y_{k}-(\hat{H}u_{k}+d_{k})\|$, the expression of $y_{k}$, and the property of the weighted norm. For controller (13), we can perform similar analysis and obtain an upper bound akin to (-A), albeit with an additional term $2\eta\|\rho_{\textup{col}}\|$ arising from $0\in\nabla\Phi_{\ell_{2},k}(u_{k}^{*})+\rho_{\textup{col}}^{\top}\partial|u_{k}|$. We incorporate the above results into (-A) and obtain (1), where the constants are
@@
In the following lemma, we characterize the property of the controller iterations (13) and (14). The objective (8) and the quadratic part of (10) are strongly convex and smooth in $u$. Let $\mu_{\Phi}$ and $L_{\Phi}$ be the corresponding constants of strong convexity and smoothness, respectively. Recall that $P$ is the matrix appearing in the weighted norm in Lemma 1.
@@
$\begin{split}\|u_{k+1}\!-\!u_{k+1}^{*}\|\leq&\alpha\|u_{k}\!-\!u_{k}^{*}\|\!+\!\eta\frac{L_{T}^{y}\|C\|}{\lambda_{\min}(P)}\lVert x_{k}\!-\!x_{\textup{ss},k}\rVert\\ &\!+\!\eta L_{T}^{y}\|H\!-\!\hat{H}\|\|u_{k}\|\!+\!\lVert u_{k+1}^{*}\!-\!u_{k}^{*}\rVert,\end{split}$ (21)
@@
$\begin{split}\lVert u_{k+1}&-u_{k+1}^{*}\rVert\stackrel{{\scriptstyle(a)}}{{\leq}}\lVert T\bigl{(}u_{k},y_{k}\bigr{)}-u_{k}^{*}\rVert+\|u_{k+1}^{*}-u_{k}^{*}\|\\ &\stackrel{{\scriptstyle(b)}}{{\leq}}\lVert T\bigl{(}u_{k},y_{k}\bigr{)}-T\bigl{(}u_{k},Hu_{k}+d_{k}\bigr{)}\rVert\\ &\quad+\lVert T\bigl{(}u_{k},Hu_{k}\!+\!d_{k}\bigr{)}\!-\!T\bigl{(}u_{k},\hat{H}u_{k}\!+\!d_{k}\bigr{)}\rVert\\ &\quad+\lVert T\bigl{(}u_{k},\hat{H}u_{k}+d_{k}\bigr{)}-u_{k}^{*}\rVert+\|u_{k+1}^{*}-u_{k}^{*}\|\\ &\stackrel{{\scriptstyle(c)}}{{\leq}}\eta L_{T}^{y}\lVert y_{k}-(Hu_{k}+d_{k})\rVert+\eta L_{T}^{y}\|H-\hat{H}\|\|u_{k}\|\\ &\quad+\alpha\|u_{k}-u_{k}^{*}\|+\lVert u_{k+1}^{*}-u_{k}^{*}\rVert\\ &\stackrel{{\scriptstyle(d)}}{{\leq}}\eta\frac{L_{T}^{y}\|C\|}{\lambda_{\min}(P)}\lVert x_{k}-x_{\textup{ss},k}\rVert+\eta L_{T}^{y}\|H-\hat{H}\|\|u_{k}\|\\ &\quad+\alpha\lVert u_{k}-u_{k}^{*}\rVert+\lVert u_{k+1}^{*}-u_{k}^{*}\rVert,\\ \end{split}$
where (a) and (b) use the triangle inequality; (c) follows from the Lipschitz continuity of $T$ and the contraction of $T$ (see [36, Proposition 25.9], where $\alpha$ is given in the lemma when $\eta\in(0,2\mu_{\Phi}/L_{\Phi}^{2})$, and we also use the non-expansiveness property of the proximal operator for (14)); and (d) applies (1) and the property of the weighted norm. ∎
@@
The main idea is to analyze the coupled evolution of state dynamics and controller iterations, whose properties are established in Lemmas 1 and 2.
@@
$\underbrace{\begin{bmatrix}\|u_{k+1}\!-\!u_{k+1}^{*}\|\\ \lVert x_{k+1}\!-\!x_{\textup{ss},k+1}\rVert_{P}\end{bmatrix}}_{\triangleq w_{k+1}}\leq\underbrace{\begin{bmatrix}\alpha&\eta\frac{L_{T}^{y}\|C\|}{\lambda_{\min}(P)}\\ \eta c_{2}&c_{1}\end{bmatrix}}_{\triangleq M}\begin{bmatrix}\|u_{k}\!-\!u_{k}^{*}\|\\ \lVert x_{k}\!-\!x_{\textup{ss},k}\rVert_{P}\end{bmatrix}$
@@
where the constants $c_{1}$ to $c_{4}$ are given by (20). Note that $M$ in (-B) is a $2$-by-$2$ positive matrix, and therefore its Perron eigenvalue equals $\rho(M)$. Hence, the requirement that $\rho(M)<1$ is equivalent to $m_{11}+m_{22}-m_{11}m_{22}+m_{21}m_{12}<1$, where $m_{ij}$ denotes the $ij$-th element of $M$. This inequality translates to
@@
where $\alpha$ and $c_{1}$ are given in Lemma 2 and (20), respectively. The function $g(\eta)$ satisfies $g(0)=1,g^{\prime}(0)<0$. Hence, there exists $\eta^{*}\in(0,2\mu_{\Phi}/L_{\Phi}^{2})$ such that for any $\eta\in(0,\eta^{*})$, $g(\eta)<1$, implying $\rho(M)<1$. We telescope (-B) and obtain
@@
When $\eta\in(0,\eta^{*})$, there exists $r>0$ and $c_{M}\in[0,1)$ such that $\lVert M^{k}\rVert\leq r(c_{M})^{k}$, see also [33, Chapter 5]. Hence, we obtain from (24) the following inequality
$\begin{split}\lVert w_{k}\rVert&\leq r(c_{M})^{k}\lVert w_{0}\rVert+\sum^{k-1}_{i=0}r(c_{M})^{k-i}\lVert q_{i+1}\rVert\\ &\stackrel{{\scriptstyle(a)}}{{\leq}}r(c_{M})^{k}\lVert w_{0}\rVert+rc_{M}\lVert\bar{q}\rVert\sum^{k-1}_{i=0}(c_{M})^{i}\\ &\stackrel{{\scriptstyle(b)}}{{\leq}}r(c_{M})^{k}\lVert w_{0}\rVert+r\frac{c_{M}}{1-c_{M}}\lVert\bar{q}\rVert,\\ \end{split}$

### v2

@@ V Numerical Experiments
where $R\in\mathbb{R}^{2n\times 2n}$ and $Q\in\mathbb{R}^{n\times n}$ are positive definite cost matrices, $\mathcal{U}_{i,k}\triangleq\{[p_{i},\,q_{i}]:0\leq p_{i}\leq p^{\text{MPP}}_{i,k},q_{\min}\leq q_{i}\leq q_{\max}\}$ is the constraint set, and $q_{\min}$ and $q_{\max}$ are lower and upper bounds on reactive power actuation, respectively.
We adopt the UNICORN 56-bus test case [36] with 25 photovoltaic inverters. Although the input-output sensitivity is a nonlinear function of $u_{k}$, we learn a constant approximation $\hat{H}$ based on power flow linearization and historical data of the injected powers and voltages. This sensitivity becomes even more inexact when the grid topology changes, specifically when the point of common coupling is switched from bus 1 to bus 26. While the uncertainty set for $\hat{H}$ is hard to characterize correctly, we tune the regularizers of (13) and (14) by gradually decreasing their values from conservative upper bounds. We augment the standard feedback optimization controller (4) and the proposed controllers (13) and (14) with projection to $\mathcal{U}_{i,k}$, use the same step size, and apply these controllers to the changed grid.
As shown in the first sub-figure of Fig. 2(a), when implemented in a new environment with sensitivity uncertainty, standard feedback optimization causes oscillations and voltage violation. Note that the dashed lines in the sub-figures on the first row denote the maximum and minimum voltage limits, which equal $1.1$ p.u. and $0.9$ p.u. (i.e., $1.1$ and $0.9$ times the base voltage), respectively. This standard controller also requires large reactive power actuation. In contrast, robust feedback optimization controllers maintain voltage stability after the point of common coupling changes. This is achieved by conservatively regulating control inputs, a consequence of regularization in the face of uncertainty. The sparsity-promoting effect of the $\ell_{1}$-regularizer is reflected in Fig. 2(c), where the reactive power injections of some inverters are zero. In comparison, the $\ell_{2}$-regularizer induces isotropic shrinkage, see Fig. 2(b). Furthermore, as shown in the sub-figures in the last row, the proposed controllers lead to less active power curtailments compared to the standard approach. Overall, robust feedback optimization effectively handles model uncertainty in this example of real-time voltage regulation.
@@ VI Conclusion
We addressed steady-state optimization of a dynamical system subject to model uncertainty by presenting robust feedback optimization, which seeks optimal closed-loop performance given all possible sensitivities falling within bounded uncertainty sets. Tractable reformulations for this min-max problem via regularized steady-state specifications were then provided. We showcased the adaptation and robustness of our controllers through theoretical tracking guarantees and numerical experiments on a distribution grid. Future avenues include tuning regularizers via differentiable programming, incorporating regularization into online sensitivity learning, and pursuing robustness against model uncertainty for nonlinear stable systems.
@@
We provide two lemmas that quantify the system dynamics (1) and the controller iterations (13) and (14).
@@
Let the conditions of Theorem 2 hold. The system dynamics (1) satisfy
@@
Recall that $x_{\textup{ss},k}=(I-A)^{-1}(Bu_{k}+d_{x,k})$ is the steady state of (1) induced by $u_{k}$ and $d_{x,k}$. Let $L_{x}^{u}\triangleq\|(I-A)^{-1}B\|$ and $L_{x}^{d}\triangleq\|(I-A)^{-1}\|$ be the Lipschitz constants of $x_{\textup{ss},k}$ with respect to $u$ and $d$, respectively. Since $\rho(A)<1$, for any given positive definite $\bar{Q}\in\mathbb{R}^{n\times n}$, there exists a unique positive definite $P\in\mathbb{R}^{n\times n}$ satisfying the Lyapunov equation $A^{\top}PA-P+\bar{Q}=0$. Therefore,
@@
$\hat{\nabla}\Phi_{\ell_{2},k}(u_{k},y_{k})$ $=2Ru_{k}\!+\!2\lambda\hat{H}^{\top}Q(y_{k}\!-\!r_{k})\!+\!2\rho_{\text{gen}}u_{k},$
@@
$\lVert u_{k+1}-u_{k}\rVert=\|u_{k}-\eta\hat{\nabla}\Phi_{\ell_{2},k}(u_{k},y_{k})-u_{k}\|$
@@
where (a) uses the triangle inequality; (b) is because of the triangle inequality and $u_{k}^{*}$ being the optimal point, i.e., $\nabla\Phi_{\ell_{2},k}(u_{k}^{*})=0$; (c) follows from the Lipschitz continuity of $\nabla\Phi_{\ell_{2},k}$, the addition and subtraction of $Hu_{k}+d_{k}$ inside $\|y_{k}-(\hat{H}u_{k}+d_{k})\|$, the expression of $y_{k}$, and the property of the weighted norm. For controller (14), we can perform similar analysis and obtain an upper bound akin to (-A), albeit with an additional term $2\eta\|\rho_{\textup{col}}\|$ because of the optimality condition $0\in\nabla\Phi_{\ell_{2},k}(u_{k}^{*})+\rho_{\textup{col}}^{\top}\partial|u_{k}|$. We incorporate the above results into (-A) and obtain (1), where the constants are given by
@@
In the following lemma, we characterize the property of the controller iterations (13) and (14). Both the objective (9) and the quadratic part of (11) are strongly convex and smooth in $u$. Let $\mu_{\Phi}$ and $L_{\Phi}$ be the corresponding constants of strong convexity and smoothness, respectively. Recall that $P$ is the matrix appearing in the weighted norm in Lemma 1.
@@
$\begin{split}\|u_{k+1}\!-\!u_{k+1}^{*}\|\leq&\alpha\|u_{k}\!-\!u_{k}^{*}\|\!+\!\eta\frac{L_{T}^{y}\|C\|}{\lambda_{\min}(P)}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lVert x_{k}\!-\!x_{\textup{ss},k}\rVert_{P}}\\ &\!+\!\eta L_{T}^{y}\|H\!-\!\hat{H}\|\|u_{k}\|\!+\!\lVert u_{k+1}^{*}\!-\!u_{k}^{*}\rVert,\end{split}$ (21)
@@
[…]
where (a) and (b) use the triangle inequality; (c) follows from the Lipschitz continuity of $T$ and the contraction of $T$ (see [37, Proposition 25.9], where $\alpha$ is given in the lemma when $\eta\in(0,2\mu_{\Phi}/L_{\Phi}^{2})$, and we also use the non-expansiveness property of the proximal operator for (14)); and (d) applies (1) and the property of the weighted norm. ∎
@@
The main idea is to analyze the coupled evolution of state dynamics and controller iterations, whose properties are established in Lemmas 1 and 2, respectively.
@@
$\underbrace{\begin{bmatrix}\|u_{k+1}\!-\!u_{k+1}^{*}\|\\ \lVert x_{k+1}\!-\!x_{\textup{ss},k+1}\rVert_{P}\end{bmatrix}}_{\triangleq w_{k+1}}\leq\underbrace{\begin{bmatrix}\alpha&\eta\frac{L_{T}^{y}\|C\|}{\lambda_{\min}(P)}\\ \eta c_{2}&c_{1}\end{bmatrix}}_{\triangleq M}\underbrace{\begin{bmatrix}\|u_{k}\!-\!u_{k}^{*}\|\\ \lVert x_{k}\!-\!x_{\textup{ss},k}\rVert_{P}\end{bmatrix}}_{\triangleq w_{k}}$
@@
where the constants $c_{1}$ to $c_{4}$ are given by (20). Note that $M$ in (-B) is a 22-by-22 positive matrix, and therefore its Perron eigenvalue equals $\rho(M)$. Hence, the requirement that $\rho(M)<1$ is equivalent to $m_{11}+m_{22}-m_{11}m_{22}+m_{21}m_{12}<1$, where $m_{ij}$ denotes the $ij$-th element of $M$. This inequality translates to
@@
where $\alpha$ and $c_{1}$ are given in Lemma 2 and (20), respectively. When $\eta=0$, we have $\alpha=1,c_{1}=\sqrt{1-\gamma}$. The function $g(\eta)$ satisfies $g(0)=1,g^{\prime}(0)=-(1-\sqrt{1-\gamma})\mu_{\Phi}<0$. Hence, there exists $\eta^{*}\in(0,2\mu_{\Phi}/L_{\Phi}^{2})$ such that for any $\eta\in(0,\eta^{*})$, $g(\eta)<1$, implying $\rho(M)<1$. We telescope (-B) and obtain
@@
When $\eta\in(0,\eta^{*})$, there exists $r>0$ and $c_{M}\in[0,1)$ such that $\lVert M^{k}\rVert\leq r(c_{M})^{k}$, see also [34, Chapter 5]. Hence, we obtain from (24) the following inequality
[…]